Jemima has three boxes whose total weight is 51 1


5 kg. Box B weighs 5 1


2 kg more than box A and box C


weighs 3 1


5 kg more than box B. Choose the correct statement.


[1] Box A weighs 17 5


6 kg.


[2] Box C weighs 21 1


30 kg.


[3] Box A weighs 42 1


2 kg.


[4] Box B weighs 12 1


3 kg.


1
Expert's answer
2022-04-15T03:33:53-0400

"x-box\,\,A\\x+5.12 -\,\,box\,\,B\\x+5.12+3.15-box\,\,C\\Then\\x+\left( x+5.12 \right) +\left( x+5.12+3.15 \right) =51.15\\3x+13.39=51.15\\3x=37.76\\x=12.5867\\x+5.12=17.7067\\x+5.12+3.15=20.8567\\No\,\,correct\,\,answer"